# Combinatorial Selection with Costly Information

### v1

@@ 9 Run the frugal Gittins policy (Singla, 2017) on the resulting mandatory-inspection instance
  2. If $K_{i}L_{i}=1$ (marked “grab”), $\mathcal{B}^{\prime}_{i}=(m_{i},0)$, i.e. the box is free to open and always contains value $m_{i}$.
Under instance $\mathbb{I}^{\prime}$, box $i$’s surrogate value is given by $W^{\mathrm{alg}}_{i}$, so
[…]
The result then follows by the law of total expectation. ∎
Lemma 7.5.
Under the hypotheses of Theorem 7.3,
$\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}W^{\mathrm{alg}}_{i}\bigg{]}\geq\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}\alpha W^{*}_{i}-\smashoperator[]{\sum_{i\in S^{\mathrm{grab}}}^{}}\beta m_{i}\bigg{]},$ (4)
Proof.
The outline of the proof is as follows. We begin with the left-hand side of 4. Then, for each box $i$, we swap $W^{\mathrm{alg}}_{i}$ with $\alpha W^{*}_{i}$, and also subtract $\beta m_{i}$ if box $i$ is marked “grab”. Because each box admits a semilocal $(\alpha,\beta)$-approximation, each of these replacements only decreases the expression’s expected value. After all $n$ replacements, we are left with the right-hand side of 4, as desired. This is the same strategy used by Scully and Doval (2024, Theorem 5.4), but some careful conditioning is needed to account for the $\beta m_{i}$ subtractions.
In order to notate the one-by-one replacement outlined above, let
$W^{(j)}_{i}$ =$=\begin{cases}W^{\mathrm{alg}}&\text{if }i\leq j\\ \alpha W^{*}&\text{if }i>j,\end{cases}$ $U^{(j)}$ $=\max_{S\in\mathcal{F}}\sum_{i\in S}W^{(j)}_{i}.$
Using this notation and recalling how $S^{\mathrm{grab}}$ is defined in Algorithm 3, we can rewrite our goal 4 as
$\mathbf{E}[U^{(n)}]\geq\mathbf{E}\bigg{[}U^{(0)}-\sum_{i=1}^{n}\beta K_{i}L_{i}m_{i}\bigg{]}.$
Therefore, it suffices to show that for all $j\in\{1,\dots,m\}$,
$\mathbf{E}[U^{(j)}]\geq\mathbf{E}\bigg{[}U^{(j-1)}-\beta K_{j}L_{j}m_{j}\bigg{]}.$ (5)
We will show 5 using Definition 11. But in order to do so, we need to express each side in terms of a maximum between $W^{\mathrm{alg}}_{j}$ or $\alpha W^{*}_{j}$ and a quantity that is independent of box $j$’s value $X_{j}$ and coin flip $K_{j}$. We express the latter quantity in terms of
[…]
These can both be seen as optimal total surrogate values achievable without box $j$. The difference is that $Y_{\neq j}$ optimizes over sets that exclude $j$, whereas $Z_{\neq j}$ optimizes over sets that include $j$ (but still excludes box $j$’s surrogate value from the sum).
With the definitions of $Y_{\neq j}$ and $Z_{\neq j}$ in hand, we can express $U^{(j)}$ and $U^{(j-1)}$ as
So to show 5, it suffices to show
$\mathbf{E}[\max\{W^{\mathrm{alg}}_{j},Y_{\neq j}-Z_{\neq j}\}]\leq\mathbf{E}[\max\{W^{\mathrm{alg}}_{j},Y_{\neq j}-Z_{\neq j}\}-\beta K_{j}L_{j}m_{j}].$
[…]
The key to showing 6 is observing the following independence facts:
The main obstacle to applying Definition 11 to 6 is that $W^{\mathrm{alg}}_{j}$ depends on $L_{j}$, which in turn depends on $K_{<j}$. Fortunately, we see from Algorithm 3 that
This is because for all $i$, when executing Algorithm 3, the only randomness the algorithm has used is the past coin flips $K_{<i}$. So to show 6, we split into cases based on whether $L_{j}=0$ or $L_{j}=1$.
[…]
So, by Definition 11,
[…]
$\geq\mathbf{E}[\max\{\alpha W^{*}_{j},y\}]-\beta p_{j}m_{j}$
$=\mathbf{E}[\max\{\alpha W^{*}_{j},y\}-\beta K_{j}m_{j}]$
[…]
The fact that $(Y_{\neq j},Z_{\neq j})$ is conditionally independent of $(K_{j},X_{j})$ given $K_{<j}$ completes the proof of 6 on the event $L_{j}=1$.
Suppose now that $L_{j}=0$. In this case, we mark box $j$ as “open” regardless of the coin flip $K_{j}$, so instead of using Definition 11, we will show that marking box $j$ as “open” does not lose any any potential value. Specifically, we will show the following:
  2. If $L_{j}=0$, then $Y_{\neq j}-Z_{\neq j}\geq m_{j}$.
Together with the fact that $m_{j}\geq h_{j}$ (1), facts (a) and (b) imply that for any $k_{<j}$ such that $K_{<j}=k_{<j}$ induces $L_{j}=0$ in Algorithm 3,
[…]
which completes the proof of 6 on the event $L_{j}=0$. It remains only to show (a) and (b). For (a), observe that if $y\geq h_{j}$, then $y_{j}=\max\{h_{j},y_{j}\}$, so
$\mathbf{E}[\max\{W^{\mathrm{open}}_{j},y\}]=\mathbf{E}[\max\{W^{\mathrm{open}}_{j},h_{j},y\}]=\mathbf{E}[\max\{W^{*}_{j},y\}].$
For (b), we will show that if $L_{j}=0$, then $Y_{\neq j}\geq m_{j}+Z_{\neq j}$. Let
  2. $S^{\mathrm{grab}}_{<j}=S^{\mathrm{grab}}\cap\{1,\dots,j-1\}$ be the boxes marked “grab” before the $i=j$ for loop iteration, and
$B^{Y}=(B^{Z}\setminus\{j\})\cup\{k\}.$
But $B^{\mathrm{grab}}\setminus B^{Z}\subseteq S^{\mathrm{grab}}_{<j}$, which means $W^{(j)}_{k}=m_{k}\geq m_{j}$. This means
[…]
7.4 Breaking the $1/2$ Barrier with Semilocal Approximation
Lemma 7.6.
Consider a box $\mathcal{B}$ satisfying 1. For all $\beta\in\mathbb{R}$ there exists a $p\in[0,1]$ that is a semilocal $(\alpha(\beta),\beta)$-approximation for $\mathcal{B}$ where
$\alpha(\beta)=\begin{cases}\frac{1}{1+\frac{c}{m}-\beta\frac{c}{m-c}}&\text{if}1\geq\frac{1}{1+\frac{c}{m}-\beta\frac{c}{m-c}}>0\\ 1&\textnormal{otherwise.}\end{cases}$
Moreover, the $p$ that achieves this is $p=\alpha(\beta)\frac{c}{m}$.
Proof.
To simplify notation we denote $\alpha(\beta)$ as $\alpha$ throughout the proof. Proving this lemma amounts to showing that this $\alpha$ and $\beta$ satisfy 3,
$(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+p\max\{m,y\}\geq\mathbf{E}[\max\{\alpha W^{*},y\}]-p\beta m$
for all $y\geq 0$. This is equivalent to showing that
$f(y)=(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+p\max\{m,y\}-\mathbf{E}[\max\{\alpha W^{*},y\}]+p\beta m$
satisfies $f(y)\geq 0$ for all $y\geq 0$. We will first show that it is sufficient to check that $f(0)\geq 0$ and that $f(y)\geq 0$ for $y\geq m$. It follows from the definition of $h$ that at all values $y$ for which $f^{\prime}(y)$ is defined, it is equal to
$(1-p)\operatorname{Pr}\left[y\geq W^{\mathrm{open}}\right]+pI(y>m)-I(y>\alpha h)\operatorname{Pr}\left[y>\alpha W^{\mathrm{open}}\right].$
Observe that
  • if $y<\alpha h<m$, then $f^{\prime}(y)\geq 0$,
  • if $\alpha h<y<m$, then since $\alpha\leq 1$, we must have $f^{\prime}(y)\leq 0$.
Thus a global minimum of $f(y)$ must be at $y=0$ or on $y\geq m$, so it is sufficient to check that 3 is satisfied when $y=0$ and $y\geq m$. When $y=0$, 3 reduces to
$(1-p)(m-c)+pm\geq\alpha m-p\beta m.$
which holds for the values of $\alpha$ and $p$ given in the lemma. When $y\geq m$, 3 reduces to
$(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+py\geq\mathbf{E}[\max\{\alpha W^{*},y\}]-p\beta m=\alpha\mathbf{E}[\max\{W^{\mathrm{open}},\frac{y}{\alpha}\}]-p\beta m$ (7)
where the last equality follows from the fact that $\mathbf{E}[\max\{\alpha W^{*},y\}]=\max\{\mathbf{E}[\max\{\alpha W^{\mathrm{open}},y\}],\alpha m\}$. The slope of $\mathbf{E}[\max\{W^{\mathrm{open}},y\}]$ is bounded above by $1$, so
$\alpha\mathbf{E}[\max\{W^{\mathrm{open}},\frac{y}{\alpha}\}]\leq\alpha\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+\alpha\left(\frac{y}{\alpha}-y\right).$
Applying this bound to 7, we find that 7 holds if,
$(\alpha+p-1)(\mathbf{E}[\max\{W^{\mathrm{open}},y\}]-y)\leq p\beta m.$ (8)
Since the slope of $\mathbf{E}[\max\{W^{\mathrm{open}},y\}]$ is bounded above by $1$, it must be that $\mathbf{E}[\max\{W^{\mathrm{open}},y\}]-y$ is maximized at $y=0$ where it is equal to $m-c$, so 8 holds if,
$(\alpha+p-1)(m-c)\leq p\beta m.$
Observe that if $p=\alpha\frac{c}{m}$, we can rewrite this inequality as
$\alpha\left(1+\frac{c}{m}-\frac{c}{m-c}\beta\right)\leq 1,$
which is satisfied by the $\alpha$ given in the lemma. ∎
Theorem 7.7.
For any box $\mathcal{B}$ satisfying 1, if $\beta=\frac{1}{10}$, the $p$ given in Lemma 7.6 is a semilocal $(\alpha,\beta)$-approximation for $\mathcal{B}$ with $\alpha\geq 0.682$.
Proof.
Lemma 7.6 guarantees that for $\beta=\frac{1}{10}$, there is some $p\in[0,1]$ that is a semilocal $(\alpha,\beta)$-approximation for $\mathcal{B}$ with
$\alpha=\begin{dcases}\frac{1}{1+\frac{c}{m}-\frac{c}{10(m-c)}}&\text{if }1\geq\frac{1}{1+\frac{c}{m}-\frac{c}{10(m-c)}}>0\\ 1&\text{otherwise.}\end{dcases}$
It is straightforward to verify (using a computer algebra system or numerical solver) that since $\frac{c}{m}\in[0,1]$, $1{\bigm{/}}\big{(}1+\frac{c}{m}-\frac{c}{10(m-c)}\big{)}\geq 0.682$ when $01\geq 1{\bigm{/}}\big{(}1+\frac{c}{m}-\frac{c}{10(m-c)}\big{)}>$ 0. ∎
Appendix
Appendix A Omitted Proofs from Section 3
@@
A.1 Amortized Surrogate Costs for Markov Chains
@@
Amortization Lower Bound.
We begin with the proof of the lower bound we stated in Lemma 3.1, relating the cost of any algorithm to the surrogate costs of any amortization. We note that one way to prove Lemma 3.1 is to follow the approach developed by Singla [2017] and relate the performance of any algorithm on a “costly information” instance to its performance in a “free information world” where action costs are paid by an outside investor who is in turn paid back the extra amortized cost of any terminal state that the algorithm accepts. Instead, we provide an algorithmic proof that will allow us to directly argue about the optimality of the water filling index policy, as well as extend our setting to MDPs.
@@
Let $\mathcal{M}_{i}=(S_{i},\sigma_{i},A_{i},c_{i},\mathcal{D}_{i},V_{i},T_{i})$ for $i\in[n]$ be the Markov chains comprising $\mathbb{I}$ and let $\mathrm{ALG}$ be any algorithm for $\mathbb{I}$. We define the following random events:
  • $I(s_{i}):=$ the event that $\mathrm{ALG}$ advances $\mathcal{M}_{i}$ at the non-terminal state $s_{i}\in S_{i}$.
  • $A(t_{i}):=$ the event that $\mathrm{ALG}$ accepts the terminal state $t_{i}\in T_{i}$.
  • $X(i):=$ the event that $\mathrm{ALG}$ accepts a terminal state of $\mathcal{M}_{i}$.
@@
Then, we can write the expected cost of the algorithm as
@@
$\begin{split}\operatorname{E}\left[\mathrm{cost}(\mathrm{ALG})\right]&=\sum_{i=1}^{n}\bigg{(}\sum_{s_{i}\in S_{i}}\operatorname{Pr}\left[I(s_{i})\right]\cdot c_{i}(s_{i})+\sum_{t_{i}\in T_{i}}\operatorname{Pr}\left[A(t_{i})\right]\cdot v_{i}(t_{i})\bigg{)}\\ &=\sum_{i=1}^{n}\bigg{(}\sum_{s_{i}\in S_{i}}\operatorname{Pr}\left[I(s_{i})\right]\cdot c_{i}(s_{i})+\sum_{t_{i}\in T_{i}}\operatorname{Pr}\left[A(t_{i})\right]\cdot\big{(}\rho_{b_{i}}(t_{i})-\sum_{s_{i}:t_{i}\in T(s_{i})}b_{s_{i}t_{i}}\big{)}\bigg{)}\\ &=\sum_{i=1}^{n}\sum_{t_{i}\in T_{i}}\operatorname{Pr}\left[A(t_{i})\right]\cdot\rho_{b_{i}}(t_{i})+\sum_{i=1}^{n}\sum_{s_{i}\in S_{i}}\bigg{(}\operatorname{Pr}\left[I(s_{i})\right]\cdot c_{i}(s_{i})-\sum_{t_{i}\in T(s_{i})}\operatorname{Pr}\left[A(t_{i})\right]\cdot b_{s_{i}t_{i}}\bigg{)}\\ \end{split}$
We will now show the following inequalities:
@@
$\sum_{i=1}^{n}\sum_{t_{i}\in T_{i}}\operatorname{Pr}\left[A(t_{i})\right]\cdot\rho_{b_{i}}(t_{i})\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$ (9)
$\operatorname{Pr}\left[I(s_{i})\right]\cdot c_{i}(s_{i})\geq\sum_{t_{i}\in T(s_{i})}\operatorname{Pr}\left[A(t_{i})\right]\cdot b_{s_{i}t_{i}}\quad\forall i\in[n],s_{i}\in S_{i}$ (10)
that will directly imply the lemma.
We begin with inequality 9. A necessary condition for the algorithm to accept a terminal state $t_{i}\in T_{i}$ is for this terminal to be realized by the Markov chain; in general, we use $p_{i}(s_{i})$ to denote the probability that a random walk on chain $\mathcal{M}_{i}$ passes through state $s_{i}\in S_{i}$ and $Q(s_{i})$ to denote the corresponding event. Then, for $A(t_{i})$ to happen a necessary condition is for $Q(t_{i})$ to happen. Furthermore, recall that by definition the surrogate cost $\rho_{i}$ corresponds to sampling a terminal state $t_{i}$ with probability $p_{i}(t_{i})$ and returning $\rho_{b_{i}}(t_{i})$. Thus:
@@
$\sum_{i=1}^{n}\sum_{t_{i}\in T_{i}}\operatorname{Pr}\left[A(t_{i})\right]\cdot\rho_{b_{i}}(t_{i})$ $=\sum_{i=1}^{n}\sum_{t_{i}\in T_{i}}\operatorname{Pr}\left[A(t_{i})|Q(t_{i})\right]\cdot p_{i}(t_{i})\cdot\rho_{b_{i}}(t_{i})$
$=\sum_{i=1}^{n}\sum_{t_{i}\in T_{i}}p_{i}(t_{i})\cdot\operatorname{E}\left[X(i)\cdot\rho_{i}|Q(t_{i})\right]$
$=\sum_{i=1}^{n}\operatorname{E}\left[X(i)\cdot\rho_{i}\right]$
$\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$
where the last inequality follows from feasibility; the set of Markov chains selected by the algorithm (i.e. the set $S=\{i:X(i)=1\}$) must be feasible, and thus $S\in\mathcal{F}$.
We now proceed to inequality 10. The proof follows from the crucial fact that for each state $s_{i}\in S_{i}$ and each terminal $t_{i}\in T_{i}(s_{i})$, it holds that $\operatorname{Pr}\left[A(t_{i})|Q(t_{i})\right]\leq\operatorname{Pr}\left[I(s_{i})|Q(s_{i})\right]$. To see why this is true, observe that for $A(t_{i})$ to happen both $Q(t_{i})$ (the algorithm cannot accept a non-realized terminal state) and $I(s_{i})$ (the algorithm needs to advance all the states in the unique path from $\sigma_{i}$ to $t_{i}$ in order to accept $t_{i}$) need to happen. Likewise, for $I(s_{i})$ to happen it must be the case that $Q(s_{i})$ happened. Chaining these events, we have
$\operatorname{Pr}\left[A(t_{i})|Q(t_{i})\right]=\frac{\operatorname{Pr}\left[A(t_{i})\right]}{p_{i}(t_{i})}=\frac{\operatorname{Pr}\left[A(t_{i})|I(s_{i})\right]\cdot\operatorname{Pr}\left[I(s_{i})|Q(s_{i})\right]\cdot p_{i}(s_{i})}{p_{i}(t_{i})}=\operatorname{Pr}\left[I(s_{i})|Q(s_{i})\right]\cdot\frac{\operatorname{Pr}\left[A(t_{i})|I(s_{i})\right]}{\operatorname{Pr}\left[Q(t_{i})|Q(s_{i})\right]}$
and the claim follows by $\operatorname{Pr}\left[A(t_{i})|I(s_{i})\right]\leq\operatorname{Pr}\left[Q(t_{i})|Q(s_{i})\right]$; this is true since even conditioned on advancing $s_{i}$, in order to accept $t_{i}$ is still needs to be realized and this happens with probability $\operatorname{Pr}\left[Q(t_{i})|Q(s_{i})\right]$. From this, we finally have that
$\sum_{t_{i}\in T(s_{i})}\operatorname{Pr}\left[A(t_{i})\right]\cdot b_{s_{i}t_{i}}$ $=\sum_{t_{i}\in T(s_{i})}\operatorname{Pr}\left[A(t_{i})|Q(t_{i})\right]\cdot p_{i}(t_{i})\cdot b_{s_{i}t_{i}}$
$\leq\sum_{t_{i}\in T(s_{i})}\operatorname{Pr}\left[I(s_{i})|Q(s_{i})\right]\cdot p_{i}(t_{i})\cdot b_{s_{i}t_{i}}$
$=\operatorname{Pr}\left[I(s_{i})|Q(s_{i})\right]\cdot\sum_{t_{i}\in T(s_{i})}p_{i}(t_{i})\cdot b_{s_{i}t_{i}}$
$=\operatorname{Pr}\left[I(s_{i})|Q(s_{i})\right]\cdot c_{i}(s_{i})\cdot p_{i}(s_{i})$
$=\operatorname{Pr}\left[I(s_{i})\right]\cdot c_{i}(s_{i}).$
where the fourth line follows by definition of amortization. This completes the proof. ∎
@@
Observe that in the proof of Lemma 3.1, we only used inequalities at two points: (i) when relating the amortized cost of the terminals that the algorithm selected to the set of terminals of minimum realized surrogate cost and (ii) when we established that $\operatorname{Pr}\left[A(t_{i})|I(t_{i})\right]\leq\operatorname{Pr}\left[Q(t_{i})|Q(s_{i})\right]$; notice that this was only required for pairs $(s_{i},t_{i})$ with $b_{s_{i}t_{i}}>0$. The first inequality can become tight if the algorithm somehow ensures that it will always accept the feasible set of realized terminal states whose total surrogate cost is minimum. The second inequality becomes tight if the algorithm can somehow ensure that whenever it advances a state $s$ sharing cost $b_{st}>0$ with one of its terminal states $t\in T(s)$, then the algorithm will always accept $t$ conditioned on it being realized. This allows us to characterize the conditions under which the lower bound is actually met by an algorithm, which we formally state as the following corollary:
Corollary A.1.
@@
Consider a matroid-min-CICS $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ over Markov chains and let $b_{i}$ be any cost amortization of $\mathcal{M}_{i}$ with surrogate cost $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ for all $i\in[n]$. Then, any algorithm that satisfies both
@@
    Surrogate Optimality. The algorithm always accepts the feasible set of terminal states of minimum total surrogate cost.

### v2

@@ Appendix C Challenges of Composing Local Conditions
We prove the claim by computing the expected cost of all possible policies for the single-selection min-CICS instance $(\mathcal{M}_{1},\mathcal{M}_{2})$ and showing that the proposed committing policy has the minimum expected cost. At the beginning, every policy needs to either take the costly action of $\mathcal{M}_{1}$ or take either action $a_{1}$ or $a_{2}$ from $\mathcal{M}_{2}$.
  • We first consider the set of policies that start by taking the costly action of $\mathcal{M}_{1}$ at a cost of $3$; if the realized terminal has value 00 then we should clearly terminate, otherwise we play the local game on $\mathcal{M}_{2}$ for $y=50$. In other words, the optimal policy that starts by taking the costly action of $\mathcal{M}_{1}$ has expected cost
    $5.53+\frac{1}{2}\cdot 0+\frac{1}{2}\cdot\min\{f_{1}(50),f_{23}(50),f_{24}(50)\}=$ 5.5 .
  • Up next, we consider the set of policies that start by taking action $a_{1}$ in $\mathcal{M}_{2}$; in that case, we should play the local game on $\mathcal{M}_{1}$ with outside option $y=5$ optimally. In other words, the optimal policy that starts by taking action $a_{1}$ on $\mathcal{M}_{2}$ has expected cost
    $\min\{5,3+\frac{1}{2}\cdot 0+\frac{1}{2}\cdot 5\}=5.$
  • Finally, we need to consider the set of policies that start by taking action $a_{2}$. If the realized outcome is the terminal state of value 00 (this happens with probability $1/2$), then we should clearly accept it and terminate at a total cost of 00. Otherwise, we are left with the task of computing the optimal policy for the single-selection min-CICS $(\mathcal{M}_{1},\mathcal{M}_{s})$ where $\mathcal{M}_{s}$ is the sub-process of $\mathcal{M}_{2}$ rooted at state $s$. In other words, the cost of the optimal policy that starts by taking action $a_{2}$ on $\mathcal{M}_{2}$ will be precisely
    $\frac{1}{2}\cdot\mathrm{OPT}(\mathcal{M}_{1},\mathcal{M}_{s}).$
    We will now compute the cost of the optimal policy for the sub-instance $(\mathcal{M}_{1},\mathcal{M}_{s})$. Observe that any policy will either start by taking $a_{3}$ or $a_{4}$ or by taking the costly action in $\mathcal{M}_{1}$; in the latter case, if the outcome is 00 the policy should terminate, otherwise there is once again an optimal choice between $a_{3}$ and $a_{4}$. Thus, we can assume without loss that the optimal policy for the sub-instance $(\mathcal{M}_{1},\mathcal{M}_{s})$ will commit in advance to either $a_{3}$ or $a_{4}$. Furthermore, since both $a_{3}$ and $a_{4}$ have a cost of 00, we can assume that we start by taking the corresponding action. Thus:
    • If the policy commits to $a_{3}$, the optimal cost for the sub-instance $(\mathcal{M}_{1},\mathcal{M}_{s})$ is
      $\min\{12,3+\frac{1}{2}\cdot 0+\frac{1}{2}\cdot\min\{12,50\}\}=9.$
    • If the policy commits to $a_{4}$, the optimal cost for the sub-instance $(\mathcal{M}_{1},\mathcal{M}_{s})$ is
      $\frac{1}{2}\cdot 0+\frac{1}{2}\cdot\min\{50,3+\frac{1}{2}\cdot\min\{0,50\}+\frac{1}{2}\cdot\min\{50,50\}\}=14.$
    Thus, the cost of the optimal policy that starts by taking action $a_{2}$ on $\mathcal{M}_{2}$ will be $(1/2)\cdot 9=4.5$.
From the above case analysis, we obtain that the optimal policy for $(\mathcal{M}_{1},\mathcal{M}_{2})$ will start by taking action $a_{2}$, then action $a_{3}$ and finally the costly action of $\mathcal{M}_{1}$ the policy terminates whenever a terminal state of value 00 is reached or it runs out of actions to take (in which case it accepts the terminal of value $12$). Clearly, this is equivalent to the optimal policy that commits to $\pi_{23}$, concluding the proof.
Appendix D Omitted Proofs from Section 3 (Amortization)
@@
D.1 Amortized Surrogate Costs for Markov Chains
@@
We begin with the proof of the lower bound we stated in Lemma 3.1, relating the cost of any algorithm to the surrogate costs of any amortization. We note that one way to prove Lemma 3.1 is to follow the approach developed by Singla (2017) and relate the performance of any algorithm on a “costly information” instance to its performance in a “free information world” where action costs are paid by an outside investor who is in turn paid back the extra amortized cost of any terminal that the algorithm follows and accepts. Instead, we provide an algorithmic proof that will allow us to directly argue about the optimality of the water filling index policy, as well as extend our setting to MDPs.
@@
Let $\mathrm{ALG}$ be any algorithm for $\mathbb{I}$. We will decompose the expected cost of $\mathrm{ALG}$ into the expected cost that it pays in every Markov chain $\mathcal{M}_{i}$ for all $i\in[n]$. Let $\mathcal{M}_{i}=(S_{i},\sigma_{i},A_{i},c_{i},\mathcal{D}_{i},V_{i},T_{i})$ be the $i$-th Markov chain and let $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ be any cost amortization for this chain. We drop the $i$-index for notational convenience and define the following random events:
  • $I(s):=$ the event that $\mathrm{ALG}$ advances $\mathcal{M}$ at the non-terminal state $s\in S$.
  • $R(\tau):=$ the event that nature realizes a trajectory $\tau^{\prime}\in\Upsilon$ with suffix $\tau$.
  • $R(s):=$ the event that nature realizes a trajectory $\tau\in\Upsilon$ with $s\in\tau$.
  • $A:=$ the event that $\mathrm{ALG}$ accepts a terminal state $t\in T$.
@@
Then, the expected cost paid by the algorithm on Markov chain $\mathcal{M}$ can be written as:
$\mathrm{ALG}(\mathcal{M})=\sum_{s\in S}\operatorname{Pr}\left[I(s)\right]\cdot c(s)+\sum_{\tau\in\Upsilon}\operatorname{Pr}\left[R(\tau)\right]\cdot\operatorname{Pr}\left[A|R(\tau)\right]\cdot v(\tau)$ (3)
For the expected cost of taking costly actions, we have that
$\sum_{s\in S}\operatorname{Pr}\left[I(s)\right]\cdot c(s)=\sum_{s\in S}\operatorname{Pr}\left[I(s)\right]\cdot\sum_{\tau\in\Upsilon(s)}\frac{\operatorname{Pr}\left[R(\tau)\right]}{\operatorname{Pr}\left[R(s)\right]}\cdot b_{s\tau}=\sum_{s\in S}\operatorname{Pr}\left[I(s)\right]\cdot\sum_{\tau\in\Upsilon(s)}\operatorname{Pr}\left[R(\tau)|R(s)\right]\cdot b_{s\tau}$ (4)
where the first equality follows by the definition of cost shares and the second equality follows by $\tau^{\prime}\in\Upsilon(s)$.
For the expected cost of accepting a terminal state, we have that
$\sum_{\tau\in\Upsilon}\operatorname{Pr}\left[R(\tau)\right]\cdot\operatorname{Pr}\left[A|R(\tau)\right]\cdot v(\tau)=\sum_{\tau\in\Upsilon}\operatorname{Pr}\left[R(\tau)\right]\cdot\operatorname{Pr}\left[A|R(\tau)\right]\cdot\big{(}\rho(\tau)-\sum_{s\in\tau}b_{s\tau_{s}}\big{)}$ (5)
by definition of the surrogate cost $\rho(\tau)$. We further decompose this expression into two terms:
  • For the first term, we have
    $\sum_{\tau\in\Upsilon}\operatorname{Pr}\left[R(\tau)\right]\cdot\operatorname{Pr}\left[A|R(\tau)\right]\cdot\rho(\tau)=\sum_{\tau\in\Upsilon}\operatorname{Pr}\left[R(\tau)\right]\cdot\operatorname{E}\left[A\cdot\rho(\tau)|R(\tau)\right]=\operatorname{E}\left[A\cdot\rho\right]$ (6)
    with the first equality following from $A\in\{0,1\}$ and the second by definition of the surrogate cost.
  • For the second term, we have
@@
    $\begin{split}\sum_{\tau\in\Upsilon}\operatorname{Pr}\left[R(\tau)\right]\cdot\operatorname{Pr}\left[A|R(\tau)\right]\cdot\sum_{s\in\tau}b_{s\tau_{s}}&=\sum_{s\in S}\sum_{\tau\in\Upsilon(s)}b_{s\tau}\cdot\sum_{\tau^{\prime}\in\Upsilon:\tau\subset\tau^{\prime}}\operatorname{Pr}\left[R(\tau^{\prime})\right]\cdot\operatorname{Pr}\left[A|R(\tau^{\prime})\right]\\ &=\sum_{s\in S}\sum_{\tau\in\Upsilon(s)}b_{s\tau}\cdot\operatorname{Pr}\left[A\cap R(\tau)\right]\end{split}$ (7)
    with the first line following by exchanging the summation order and the second line following by definition of the event $R(\tau)$, specifically that $R(\tau)=\sum_{\tau^{\prime}\in\Upsilon:\tau\subset\tau^{\prime}}R(\tau^{\prime})$.
Combining Equations (3) through (7), we obtain that the expected cost of the algorithm on Markov chain $\mathcal{M}$ can be written as
@@
$\mathrm{ALG}(\mathcal{M})=\operatorname{E}\left[A\cdot\rho\right]+\sum_{s\in S}\sum_{\tau\in\Upsilon(s)}b_{s\tau}\cdot\bigg{(}\operatorname{Pr}\left[I(s)\right]\cdot\operatorname{Pr}\left[R(\tau)|R(s)\right]-\operatorname{Pr}\left[A\cap R(\tau)\right]\bigg{)}.$
Observe that the second term is always non-negative: in order to accept a terminal from the chain and for nature to realize a trajectory $\tau\in\rho(s)$, it is necessary for the algorithm to advance state $s$ (this already requires that $s$ belongs in the realized trajectory) and for the suffix $\tau$ to be realized conditioned on having reached $s$; notice that the last two events are independent. Thus, we have concluded that
$\mathrm{ALG}(\mathcal{M})\geq\operatorname{E}\left[A\cdot\rho\right].$
The proof is completed by noting that
@@
$\operatorname{E}\left[\text{cost}(\mathrm{ALG})\right]=\sum_{i=1}^{n}\mathrm{ALG}(\mathcal{M}_{i})\geq\sum_{i=1}^{n}\operatorname{E}\left[A_{i}\cdot\rho_{i}\right]=\operatorname{E}\left[\sum_{i=1}^{n}A_{i}\rho_{i}\right]\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$
with the last inequality obtained by the fact that the set $S=\{i\in[n]:A_{i}=1\}$ must be feasible (i.e. $S\in\mathcal{F}$) with probability $1$ for $\mathrm{ALG}$ to be a valid policy. ∎
@@
Observe that in the proof of Lemma 3.1, we only used inequalities at two points: (i) when relating the amortized cost of the trajectories that the algorithm selected to the (feasible) set of trajectories of minimum total surrogate cost and (ii) when we established that $\operatorname{Pr}\left[I(s)\right]\cdot\operatorname{Pr}\left[R(\tau)|R(s)\right]\geq\operatorname{Pr}\left[A\cap R(\tau)\right]$ for all states $s$ and all trajectories $\tau\in\Upsilon(s)$; notice that this was only required for pairs $(s,\tau)$ with $b_{s\tau}>0$. The first inequality can become tight if the algorithm somehow ensures that it will always accept the feasible set of realized trajectories whose total surrogate cost is minimum. The second inequality becomes tight if the algorithm can somehow ensure that whenever it advances a state $s$ sharing cost $b_{s\tau}>0$ with one of its downstream trajectories $\tau$, it will always accept the corresponding Markov chain if the suffix $\tau$ gets realized. This allows us to characterize the conditions under which the lower bound is actually met by an algorithm, which we formally state as the following corollary:
Corollary D.1.
@@
Consider any instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-min-CICS over Markov chains and let $b_{i}$ be any cost amortization of $\mathcal{M}_{i}$ with surrogate cost $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ for all $i\in[n]$. Then, any algorithm that satisfies both
@@
    Surrogate Optimality. The algorithm always accepts a feasible set of Markov chains whose realized trajectories have the minimum total surrogate cost.